An economist reports that 182 of a random sample of 400 middle-income Americans actively participate in the stock market.

Formulate the 95% confidence interval for the proportion of middle-income Americans who actively participate in the stock market.

You are filling in this template:

x = (y)(z). That is, what is x?y?z?
Answer #1

The 95% confidence interval of the proportion of middle-income Americans who actively participate in the stock market = [, ], where, = 182/400 = 0.455, n = 400

Thus, 95% C.I. = [0.455 - 0.0488, 0.455 + 0.0488] = [0.4062, 0.5038] (Ans).